OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation quantity x minus five squared divided by eighty one minus the quantity of y minus one squared divided by one hundred and forty four = 1

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

waddup?

OpenStudy (anonymous):

\[\frac{(x-5)^2}{81}-\frac{(y-1)^2}{4}=1\]?

OpenStudy (anonymous):

this is way easier than you think but you need to know two things first a) the center and b) which way it is oriented

OpenStudy (anonymous):

I know that the hyperbolas are oriented left/right

OpenStudy (anonymous):

i thought ya all that earlier

OpenStudy (anonymous):

taught*

OpenStudy (anonymous):

center?

OpenStudy (anonymous):

@satellite73 you mean what i taught him?

OpenStudy (anonymous):

(0,5) ?

OpenStudy (anonymous):

My bad, (5, 1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so now what can I do?

OpenStudy (anonymous):

find c

OpenStudy (anonymous):

\[\frac{(x-5)^2}{81}-\frac{(y-1)^2}{4}=1\]\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] what is \(a\)?

OpenStudy (anonymous):

Ok. c = 15

OpenStudy (anonymous):

a = 81, b =144

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[a^2=81,b^2=144\]

OpenStudy (anonymous):

\[c^2=a^2+b^2\]

OpenStudy (anonymous):

thats why i told ya to put em into ^2 form

OpenStudy (anonymous):

Oh sorry. a = 9, b = 12

OpenStudy (anonymous):

Ok, so c = 15 because c^2 = 225

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

you know the hyperbola is transverse in x-axis so ( center +/- c, center) is your foci

OpenStudy (anonymous):

and same for the vertices

OpenStudy (anonymous):

Wait. These are my choices: A. Vertices: (17, 1), (-7, 1); Foci: (-7, 1), (17, 1) B. Vertices: (14, 1), (-4, 1); Foci: (-10, 1), (20, 1) C. Vertices: (1, 14), (1, -4); Foci: (1, -10), (1, 20) D. Vertices: (1, 17), (1, -7); Foci: (1, -7), (1, 17) So I know it has to be between A and B.

OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±3) and foci at (0, ±7). If my answer choices are: A. y squared over nine minus x squared over forty = 1 B. y squared over forty minus x squared over nine = 1 C. y squared over forty nine minus x squared over nine = 1 D. y squared over nine minus x squared over forty nine = 1 I was thinking it was between B and D.

OpenStudy (anonymous):

lemme check

OpenStudy (anonymous):

you see foci is c^2 = a^2+b^2 so we get 7^2 - 3^2 = 40

OpenStudy (anonymous):

Oh ok. So that would make it A.